# MOOSE-Chem: Large Language Models for Rediscovering Unseen Chemistry Scientific Hypotheses

### v4

@@ Abstract
Scientific discovery contributes largely to human society’s prosperity, and recent progress shows that LLMs could potentially catalyze this process. However, it is still unclear whether LLMs can discover novel and valid hypotheses in chemistry. In this work, we investigate this central research question: Can LLMs automatically discover novel and valid chemistry research hypotheses given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question? After extensive discussions with chemistry experts, we propose an assumption that a majority of chemistry hypotheses can result from a research background and several inspirations. With this key insight, we break the central question into three smaller fundamental questions. In brief, they are: (1) given a background question, whether LLMs can retrieve good inspirations; (2) with background and inspirations, whether LLMs can lead to hypothesis; and (3) whether LLMs can identify good hypotheses to rank them higher. To investigate these questions, we construct a benchmark consisting of 51 chemistry papers published in Nature, Science, or a similar level in 2024 (all papers are only available online since 2024). Every paper is divided by chemistry PhD students into three components: background, inspirations, and hypothesis. The goal is to rediscover the hypothesis, given only the background and a large randomly selected chemistry literature corpus consisting of the ground truth inspiration papers, with LLMs trained with data up to 2023. We also develop an LLM-based multi-agent framework  [^1] Code and Benchmark are available at https://github.com/ZonglinY/MOOSE-Chem.git that leverages the assumption, consisting of three stages reflecting the three smaller questions. The proposed method can rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations. Contribution during internship at Shanghai Artificial Intelligence Laboratory. Corresponding author.
@@ 1 Introduction
Yang et al. (2024b) first find that LLMs can generate novel and valid enough hypotheses evaluated by experts. They focus on the social science domain and make discoveries by developing a multi-agent system, leveraging an assumption that a majority of social science hypotheses can be divided into a research background concept and an inspiration concept. This assumption is largely valid because a social science hypothesis is about how an independent variable can influence another dependent variable (Hair et al., 2007). Si et al. (2024) further validate this finding by employing a large group of scientists to evaluate LLMs’ generated hypotheses in the NLP domain and show that LLM can generate more novel but slightly less valid research hypotheses than human researchers. However, it is still unclear LLMs’ scientific discovery ability in natural science such as the chemistry domain.
@@ 3 Benchmark Construction
The reason we add a strict version for background question and background survey is that many hypotheses are making relatively minor modifications based on existing methods covered by the survey, and the question can be very insightful to provide a hint on the general direction of the hypothesis. In practice, these situations are entirely possible, especially when the scientist users can provide a more comprehensive survey on existing methods, or contain deep insights in their question. Here we also keep the strict version to make the task more challenging and encourage developing methods to better assist scientists even when they are also new to their research topic.
@@ 4.1 Fundamental Assumption and Following Decomposition
Here $k\in\mathbb{Z}$ represents the number of inspirations needed for a particular $h$. Typically in chemistry, $k\in[1,3]$. In other words, given existing knowledge in the background, a majority of chemistry research is about searching knowledge that previously not known to be related to the background but in fact can assist the background, then associate the background knowledge and the searched knowledge in a reasonable way to compose a hypothesis.
Based on this assumption, we can transform the seemingly impossible-to-solve $P(h|b)$ into an equivalent form, where each step in the equivalent form is practical and executable.
$P(h|b)\approx\prod_{j=1}^{k}P(i_{j}|b,h_{j-1},I)\cdot P(h_{j}|b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset$ (2)
Here $I$ denotes the full (chemistry) literature, representing the full inspiration space to search for every single $i$. The full proof along with detailed analyses is shown in § A.1. Equation 2 is meaningful in that by decomposing $P(h|b)$ into more practical and executable smaller questions, the seemingly impossible-to-solve $P(h|b)$ itself becomes practical. We analyse how $P(i_{j}|b,h_{j-1},I)$ and $P(h_{j}|b,i_{j},h_{j-1})$ are practical and executable by LLMs in § 5.1 and § 5.2 correspondingly.
Now it is clear on the steps to obtain $h$ from $b$. But it still might not be enough helpful in practice, since $I$ can be on a large scale, and the search process might find lots of $i$, and finally lead to lots of $h$. Moreover, it is very time-consuming for scientists to conduct experiments to verify every single $h$. Therefore, it would be very helpful if the generated $h$ could be ranked based on quality.
Here we adopt a straightforward and efficient way for ranking. Specifically, we design a rating function $R(h)$, such that $R(h)\rightarrow\mathbb{R}$. Denoting the full set of generated $h$ as $H$, we can obtain
@@
Supported by Equation 2 and Equation 3, as a result, to model $P(h|b)$, the only three components we need to model are $P(i_{j}|b,h_{j-1},I)$, $P(h_{j}|b,i_{j},h_{j-1})$, and $R(h)$. The implementation details of the three components are illustrated in the remaining subsections in § 4. Analyses of LLM’s ability on the three components are provided in § 5.
@@ 4.2.1 The General Picture
Our methodology is developed based on the fundamental assumption discussed in § 4.1. Specifically, we use LLMs to perform $P(i_{j}|b,h_{j-1},I)$, $P(h_{j}|b,i_{j},h_{j-1})$, and $R(h)$, and organize them into a multi-agent LLM-based framework. The input to the framework is only a background question and/or background survey, together with a (large) chemistry literature corpus to search for inspiration. The output of the framework is a list of ranked research hypothesis.
@@
In the general picture, given a research background $b$ (research question and/or research survey), the framework first performs $P(i_{1}|b,h_{0}=\emptyset,I)$ by screening through the literature corpus $I$ to select many papers $i$, where each of them has the potential to serve as an inspiration. Then the framework performs $P(h_{1}|b,i_{1},h_{0}=\emptyset)$, associating $b$ and each $i$ together to compose $h$. Then, it ranks $h$ by assigning an evaluation score $r$ on each of $h_{1}$ by $R(h_{1})$. We call these three steps as one round. Another round means going through the three steps again, based on the previous round’s results.
@@
4.2.2 Design Details of $P(i_{j}|b,h_{j-1},I)$ And Its Motivation
We use LLMs to conduct a screening process for $P(i_{j}|b,h_{j-1},I)$. Specifically, for each inference, we (1) sequentially select a fixed number of papers from $I$, where the fixed number is called the screening window size (default is 15); (2) set up a prompt consisting of $b$, the title and abstract of the selected papers from $I$, and the previous $h$ (if it is not $\emptyset$); and (3) instruct the LLM to generate three titles from the input that can best serve as $i$ for $b$ (and optionally previous $h$), and give reasons.
Here we use LLMs to choose potential inspiration $i$, but not choose $i$ from citation nor semantic neighbors because $i$ is supposed to be previously not known to be related to $b$ (we have discussed it in § 4.1). If the chosen $i$ is already known to be related to $b$, then the composed $h$ probably would not be novel. If the chosen $i$ contains similar semantic information with $b$, then probably it is not necessary to add $i$ at all, since it does not introduce much (any) extra information.
Here we have a bold assumption that the most advanced LLMs, after training on hundreds of millions of scientific literature, might already know many knowledge pairs that can be associated to create novel knowledge, where the knowledge pairs are not known by any scientist to be related. It might not be too bold, since Tshitoyan et al. (2019) have shown that word embedding obtained from unsupervised learning on 3.3 million material science publication abstracts can recommend materials for functional applications several years before their discovery. Here, the functional applications can be seen as $b$, and the recommended materials can be seen as $i$, or even directly as $h$ if it is enough similar. It probably indicates that LLMs trained with significantly more literature tokens and significantly more parameters might already be able to identify the relation between many knowledge pairs that are unknown to be related by any scientist. We analyze this assumption in § 5.1.
4.2.3 Design Details of $P(h_{j}|b,i_{j},h_{j-1})$ And Its Motivation
@@
Specifically, given $b$ and $i$, EU will first generate multiple hypothesis “mutations” $m$, where each $m$ is a unique way to associate $b$ and $i$ together. Then EU further develops each $m$ independently by providing feedback to each $m$ in terms of validness, novelty, clarity, and significance, and then refining them based on the feedback. Yang et al. (2024b) first propose to provide feedback in terms of validness, novelty, and clarity to refine hypotheses. Here we add an additional aspect, significance, since significance is an important evaluation criterion in chemistry. We assume the refined hypothesis should be of better quality so that the refined hypothesis is “selected”, while the previous hypothesis is “eliminated” by the “environment”. Finally EU “recombines” the remaining selected $m$, leveraging the advantages from every $m$ to propose $h$ to better associate $b$ and $i$.
@@ 4.2.4 Design Details of $R(h)$ And Its Motivation
We adopt a simple and efficient way for $R(h)$, which is to prompt an LLM to output evaluation scores for an input $h$ in terms of validness, novelty, significance, and potential. Validness and novelty are two fundamental requirements for such an inductive reasoning process as scientific discovery (Yang et al., 2024a; b). Significance is added because it is important for chemistry. We additionally add potential, because the generated $h$ are about to be further developed by scientists, so we might want to pick those $h$ that not only are currently in high quality but also have good potential to be further developed. We didn’t design $R(h)$ in a more complicated way, since there are lots of $h$ to rank, and we might want to save more inference time.
@@ 5 Investigation on Fundamental Questions
$P(h|b)$ can be understood as the task to discover high-quality chemistry research hypothesis, given only a background question and/or background survey. Our central question to investigate is how well LLMs can perform $P(h|b)$. Supported by Equation 2 and 3, we break up this main question into three smaller questions: how well can LLMs perform (1) $P(i_{j}|b,h_{j-1},I)$, (2) $P(h_{j}|b,i_{j},h_{j-1})$, and (3) $R(h)$? All experiments are performed by GPT-4o (its training data is up to October 2023).
5.1 How Well can LLMs perform $P(i_{j}|b,h_{j-1},I)$?
Here we investigate the question (denoted as $Q1$): “whether LLM can identify inspiration papers which are unknown to be able to associate with the background (or at least unknown to associate in a certain way) but in fact can associate with the background to create novel knowledge?”.
@@
5.2 How Well can LLMs perform $P(h_{j}|b,i_{j},h_{j-1})$?
@@
Here we investigate the question (denoted as $Q2$): “Given only known knowledge, whether LLM can reason to unknown knowledge that has high probability to be valid?”.
@@ 5.3 How Well can LLMs perform $R(h)$?
Here we investigate $Q3$: “whether LLMs can select high-quality $h$ to rank them higher?”.
@@ 6 Experiment and Ablation Study
Here, we perform experiments in a setting similar to the copilot in the wild setting. Specifically, only background question (strict), background survey (strict), and a chemistry corpus $|I|=300$ are provided to the framework. Only the top 4% of $I$ is selected and used to develop $h$. The evaluation metrics are Top MS and Average MS (the highest/average Matched Score of all generated $h$ from one $b$), averaging across the benchmark. All experiments are conducted by GPT-4o (its training data is up to October 2023).
@@ 6.1 Baselines
is a hypothesis discovery framework for the NLP and biochemical domain. It relies on semantic and citation neighbors to retrieve information to assist the background. As a result, the retrieved information could be very related to the background that might not be able to serve as an inspiration. To make the generated hypothesis more novel, it adopts self-refine to focus on improving the novelty aspect of the generated hypothesis. Here we implement SciMON with LLM-based inspiration retrieval, the same as MOOSE-Chem. Table 3 shows that the recall rate of LLM-based retrieval is 83.7%.
@@ 6.2 Results
To better understand the performance of MOOSE-Chem in this real copilot setting, for each $b$ the top 4 generated $h$ with the highest MS by GPT-4o are evaluated again by two experts in terms of MS. Table 11 shows the expert evaluation results. Here the top MS is the highest MS for each $b$, out of the 4 expert evaluated $h$ for this $b$. Note that MS rated as three is already very high. Illustrated in Table 6, it means the generated $h$ by MOOSE-Chem (that has not seen $h$) in the real copilot setting covers two main innovations of the chemistry hypothesis, which is published in Nature, Science or a similar level. Some case studies can be seen in § A.16.
@@ 7 Conclusion
We investigate this central question: “Can LLMs automatically discover novel and valid chemistry (including material science) research hypotheses (even those which deserve a publication in Nature, Science, or a similar level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We propose a fundamental assumption to break up this seemingly impossible-to-solve central question into three smaller, more practical, and executable fundamental questions. Then, we investigate LLM’s ability on each of them. To do it, we construct a benchmark consisting of chemistry and material science papers published and only be public in 2024. We also develop an LLM-based multi-agent framework consisting of three stages reflecting the three smaller fundamental questions. Experiments show that the framework (runs in a copilot in-the-wild setting, with LLMs with training data up to October 2023) can rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.

### v5

@@ Abstract
Scientific discovery plays a pivotal role in advancing human society, and recent progress in large language models (LLMs) suggests their potential to accelerate this process. However, it remains unclear whether LLMs can autonomously generate novel and valid hypotheses in chemistry. In this work, we investigate whether LLMs can discover high-quality chemistry hypotheses given only a research background—comprising a question and/or a survey—without restriction on the domain of the question. We begin with the observation that hypothesis discovery is a seemingly intractable task. To address this, we propose a formal mathematical decomposition grounded in a fundamental assumption: that most chemistry hypotheses can be composed from a research background and a set of inspirations. This decomposition leads to three practical subtasks—retrieving inspirations, composing hypotheses with inspirations, and ranking hypotheses—which together constitute a sufficient set of subtasks for the overall scientific discovery task. We further develop an agentic LLM framework, MOOSE-Chem, that is a direct implementation of this mathematical decomposition. To evaluate this framework, we construct a benchmark of 51 high-impact chemistry papers published and online after January 2024, each manually annotated by PhD chemists with background, inspirations, and hypothesis. The framework is able to rediscover many hypotheses with high similarity to the groundtruth, successfully capturing the core innovations—while ensuring no data contamination since it uses an LLM with knowledge cutoff date prior to 2024. Finally, based on LLM’s surprisingly high accuracy on inspiration retrieval, a task with inherently out-of-distribution nature, we propose a bold assumption: that LLMs may already encode latent scientific knowledge associations not yet recognized by humans. Contribution during internship at Shanghai Artificial Intelligence Laboratory. Corresponding author.
@@ 1 Introduction
Yang et al. (2024b) first find that LLMs can generate novel and valid enough hypotheses evaluated by experts. They focus on the social science domain and make discoveries by developing a multi-agent system, leveraging an assumption that a majority of social science hypotheses can be divided into a research background concept and an inspiration concept. This assumption is largely valid because a social science hypothesis is about how an independent variable can influence another dependent variable (Hair et al., 2007).
Si et al. (2024) further validate this finding by employing a large group of scientists to evaluate LLMs’ generated hypotheses in the NLP domain and show that LLM can generate more novel but slightly less valid research hypotheses than human researchers. However, it is still unclear LLMs’ scientific discovery ability in natural science such as the chemistry domain.
@@ 3 Benchmark Construction
The reason we add a strict version for background question and background survey is that many hypotheses are making relatively minor modifications based on existing methods covered by the survey, and the question can be very insightful to provide a hint on the general direction of the hypothesis. In practice, these situations are entirely possible, especially when the scientist users can provide a more comprehensive survey on existing methods, or contain deep insights in their question. Here, we also keep the strict version to make the task more challenging and encourage developing methods to better assist scientists even when they are also new to their research topic.
@@ 4.1 Fundamental Assumption and Following Decomposition
Here, $k\in\mathbb{Z}$ represents the number of inspirations needed for a particular $h$. Typically in chemistry, $k\in[1,3]$. In other words, given existing knowledge in the background, a majority of chemistry research is about searching knowledge that previously not known to be related to the background but in fact can assist the background, then associate the background knowledge and the searched knowledge in a reasonable way to compose a hypothesis.
Based on this assumption, we can transform the seemingly impossible-to-solve $P(h\mid b)$ into an equivalent form, where each step in the equivalent form is practical and executable.
$P(h\mid b)\approx\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset$ (2)
Here, $I$ denotes the full (chemistry) literature, representing the full inspiration space to search for every single $i$. The full proof along with detailed analyses is shown in § A.1, which is the core of this paper (and therefore highly recommend to take a read).
Equation 2 is meaningful in that by decomposing $P(h\mid b)$ into more practical and executable smaller questions, the seemingly impossible-to-solve $P(h\mid b)$ itself becomes practical. We analyze how $P(i_{j}\mid b,h_{j-1},I)$ and $P(h_{j}\mid b,i_{j},h_{j-1})$ are practical and executable by LLMs in § 5.1 and § 5.2 correspondingly.
Now we have clarified the steps to obtain $h$ from $b$. However, it still might not be enough helpful in practice, since $I$ can be on a large scale, and the search process might find lots of $i$, and finally lead to lots of $h$. Moreover, it is very time-consuming for scientists to conduct experiments to verify every single $h$. Therefore, it would be very helpful if the generated $h$ could be ranked based on quality. Here, we adopt a straightforward and efficient way for ranking. Specifically, we design a rating function $R(h)$, such that $R(h)\rightarrow\mathbb{R}$. Denoting the full set of generated $h$ as $H$, we can obtain
@@
Supported by Equation 2 and Equation 3, as a result, to model $P(h\mid b)$, the only three components we need to model are $P(i_{j}\mid b,h_{j-1},I)$, $P(h_{j}\mid b,i_{j},h_{j-1})$, and $R(h)$. The implementation details of the three components are illustrated in the remaining subsections in § 4. Analyses of LLM’s ability on the three components are provided in § 5.
@@ 4.2.1 The General Picture
Our methodology is developed based on the fundamental assumption discussed in § 4.1. Specifically, we use LLMs to perform $P(i_{j}\mid b,h_{j-1},I)$, $P(h_{j}\mid b,i_{j},h_{j-1})$, and $R(h)$, and organize them into a multi-agent LLM-based framework. The input to the framework is only a background question and/or background survey, together with a (large) chemistry literature corpus to search for inspiration. The output of the framework is a list of ranked research hypothesis.
@@
In the general picture, given a research background $b$ (research question and/or research survey), the framework first performs $P(i_{1}\mid b,h_{0}=\emptyset,I)$ by screening through the literature corpus $I$ to select many papers $i$, where each of them has the potential to serve as an inspiration. Then the framework performs $P(h_{1}\mid b,i_{1},h_{0}=\emptyset)$, associating $b$ and each $i$ together to compose $h$. Then, it ranks $h$ by assigning an evaluation score $r$ on each of $h_{1}$ by $R(h_{1})$. We call these three steps as one round. Another round means going through the three steps again, based on the previous round’s results.
@@
4.2.2 Design Details of $P(i_{j}\mid b,h_{j-1},I)$ And Its Motivation
We use LLMs to conduct a screening process for $P(i_{j}\mid b,h_{j-1},I)$. Specifically, for each inference, we (1) sequentially select a fixed number of papers from $I$, where the fixed number is called the screening window size (default is 15); (2) set up a prompt consisting of $b$, the title and abstract of the selected papers from $I$, and the previous $h$ (if it is not $\emptyset$); and (3) instruct the LLM to generate three titles from the input that can best serve as $i$ for $b$ (and optionally previous $h$), and give reasons.
In particular, we use LLMs to choose potential inspiration $i$, but not choose $i$ from citation nor semantic neighbors because $i$ is supposed to be previously not known to be related to $b$ (we have discussed it in § 4.1). If the chosen $i$ is already known to be related to $b$, then the composed $h$ probably would not be novel. If the chosen $i$ contains similar semantic information with $b$, then probably it is not necessary to add $i$ at all, since it does not introduce much (any) extra information.
Our bold assumption here is that advanced LLMs, trained on vast scientific literature, may already recognize novel knowledge pairs unknown to any scientist that can be associated to create novel knowledge. However, this may not be too bold, as Tshitoyan et al. (2019) showed that unsupervised word embeddings from 3.3 million materials science abstracts could predict functional materials years before their discovery. Here, the functional applications can be seen as $b$, and the recommended materials can be seen as $i$, or even directly as $h$ if it is enough similar. It probably indicates that LLMs trained with significantly more literature tokens and parameters might already be able to identify the relation between many knowledge pairs that are unknown to be related by any scientist. We analyze this assumption in § 5.1.
4.2.3 Design Details of $P(h_{j}\mid b,i_{j},h_{j-1})$ And Its Motivation
@@
Specifically, given $b$ and $i$, EU will first generate multiple hypothesis “mutations” $m$, where each $m$ is a unique way to associate $b$ and $i$ together. Then EU further develops each $m$ independently by providing feedback to each $m$ in terms of validness, novelty, clarity, and significance, and then refining them based on the feedback. Yang et al. (2024b) first propose to provide feedback in terms of validness, novelty, and clarity to refine hypotheses. Here, we add an additional aspect, significance, since significance is an important evaluation criterion in chemistry. We assume the refined hypothesis should be of better quality so that the refined hypothesis is “selected”, while the previous hypothesis is “eliminated” by the “environment”. Finally EU “recombines” the remaining selected $m$, leveraging the advantages from every $m$ to propose $h$ to better associate $b$ and $i$.
@@ 4.2.4 Design Details of $R(h)$ And Its Motivation
We adopt a simple and efficient way for $R(h)$, which is to prompt an LLM to output evaluation scores for an input $h$ in terms of validness, novelty, significance, and potential. Validness and novelty are two fundamental requirements for such an inductive reasoning process as scientific discovery (Yang et al., 2024a; b). Significance is added because it is important for chemistry. We additionally add potential, because the generated $h$ are about to be further developed by scientists, so we might want to pick those $h$ that not only are currently in high quality but also have good potential to be further developed. We did not design $R(h)$ in a more complicated way, since there are lots of $h$ to rank, and we might want to save more inference time.
@@ 5 Investigation on Fundamental Questions
$P(h\mid b)$ can be understood as the task to discover high-quality chemistry research hypothesis, given only a background question and/or background survey. Our central question to investigate is how well LLMs can perform $P(h\mid b)$. Supported by Equation 2 and 3, we break up this main question into three smaller questions: how well can LLMs perform (1) $P(i_{j}\mid b,h_{j-1},I)$, (2) $P(h_{j}\mid b,i_{j},h_{j-1})$, and (3) $R(h)$? All experiments are performed by GPT-4o (its training data is up to October 2023).
5.1 How Well can LLMs perform $P(i_{j}\mid b,h_{j-1},I)$?
Here, we investigate the question (denoted as $Q1$): “whether LLM can identify inspiration papers which are unknown to be able to associate with the background (or at least unknown to associate in a certain way) but in fact can associate with the background to create novel knowledge?”.
@@ 5.1 How Well can LLMs perform $P(i_{j}\mid b,h_{j-1},I)$?
5.2 How Well can LLMs perform $P(h_{j}\mid b,i_{j},h_{j-1})$?
@@
Here, we investigate the question (denoted as $Q2$): “Given only known knowledge, whether LLM can reason to unknown knowledge that has high probability to be valid?”.
@@ 5.3 How Well can LLMs perform $R(h)$?
Here, we investigate $Q3$: “whether LLMs can select high-quality $h$ to rank them higher?”.
@@ 6 Experiment and Ablation Study
We perform experiments in a setting similar to the copilot in the wild setting. Only background question (strict), background survey (strict), and a chemistry corpus $|I|=300$ are provided to the framework. Only the top 4% of $I$ is selected and used to develop $h$. The evaluation metrics are Top MS and Average MS (the highest/average Matched Score of all generated $h$ from one $b$), averaging across the benchmark. Experiments are conducted by GPT-4o (training data up to October 2023).
@@ 6.1 Baselines
is a hypothesis discovery framework for the NLP and biochemical domain. It relies on semantic and citation neighbors to retrieve information to assist the background. As a result, the retrieved information could be very related to the background that might not be able to serve as an inspiration. To make the generated hypothesis more novel, it adopts self-refine to focus on improving the novelty aspect of the generated hypothesis. Here, we implement SciMON with LLM-based inspiration retrieval, the same as MOOSE-Chem. Table 3 shows that the recall rate of LLM-based retrieval is 83.7%.
@@ 6.2 Results
To better understand the performance of MOOSE-Chem in this real copilot setting, for each $b$ the top 4 generated $h$ with the highest MS by GPT-4o are evaluated again by two experts in terms of MS. Table 11 shows the expert evaluation results. Here, the top MS is the highest MS for each $b$, out of the 4 expert evaluated $h$ for this $b$. Note that MS rated as three is already very high. Illustrated in Table 6, it means the generated $h$ by MOOSE-Chem (that has not seen $h$) in the real copilot setting covers two main innovations of the chemistry hypothesis, which is published in Nature, Science or a similar level. Some case studies can be seen in § A.16.
@@ 7 Conclusion
We investigated this central question: “Can LLMs automatically discover novel and valid chemistry (including material science) research hypotheses (even those which deserve a publication in Nature, Science, or a similar level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We proposed a fundamental assumption to break up this seemingly impossible-to-solve central question into three smaller, more practical, and executable fundamental questions. Then, we investigated LLM’s ability on each of them.
To this end, we constructed a benchmark consisting of chemistry and material science papers published and only be public in 2024. We also developed an LLM-based multi-agent framework consisting of three stages reflecting the three smaller fundamental questions. Experiments showed that the framework (runs in a copilot in-the-wild setting, with LLMs with training data up to October 2023) can rediscover many hypotheses with very high similarity with the ground-truth ones, covering the main innovations.